# cmaes : A Simple yet Practical Python Library for CMA-ES

### v2

@@ 3.4. Unexpected Error Identification via Fuzzing
Fuzzing is an automated software testing method aimed at discovering unexpected inputs that could cause program crashes. It involves generating a wide range of inputs. The key component in this process is the fuzzing engine, which automatically generates various inputs for the target program. A notable technique is coverage-guided fuzzing. This approach not only generates random inputs but also collects code coverage data. When an input is found to increase the code coverage, indicating untested paths, this information is used to methodically explore different branches within the program’s code. This strategy enhances the detection of unexpected errors, significantly contributing to the reliability of the software.
@@ 3.6. Integration with Real-World Systems:
A key factor in Optuna’s integration of cmaes is not only its clean code and simple design but also the reduced size of its pickle serialization. In hyperparameter optimization, the focus of Optuna, unexpected errors may occur during the training of machine learning models. Consequently, it is crucial to make the optimization process resumable, necessitating the regular saving of CMA-ES optimization data, including distribution parameters. Therefore, reducing the pickle size is critical for Optuna. Our library has introduced specialized __getstate__ and __setstate__ methods that are invoked during pickle serialization and deserialization. These modifications have successfully achieved further lightweighting of the pickle objects. Figure 5 shows the significant reduction of the serialization size for the CMA-ES objects using pickle in cmaes, compared to pycma.
Figure 5. Comparison of the serialization size for the CMA-ES objects using pickle in pycma and cmaes.
@@ 4. Examples with Recent Advances
  • LRA-CMA (presented at GECCO’23 and nominated for the best paper in the ENUM Track) (Nomura et al., 2023): LRA-CMA can solve multimodal and noisy problems without expensive hyperparameter tuning.
  • WS-CMA (presented at AAAI’21) (Nomura et al., 2021): WS-CMA can solve a current task efficiently by utilizing results of a similar task and setting good initial distributions (i.e. transfer learning).
  • CMAwM (presented at GECCO’22 and nominated for the best paper in the ENUM Track) (Hamano et al., 2022): CMAwM can stably solve mixed-integer problems by modifying distribution parameters so that generated samples are not fixed.
Importantly, these methods are a generalization of CMA-ES, which means that they do not change the original implementation of CMA-ES. This enables the preservation of the library’s simplicity, making it easier for those who initially learned CMA-ES to straightforwardly delve into understanding these methods.
@@ 4.2. Transfer Learning
Why do we need WS-CMA?.
Practically it is possible that we have already solved a problem (source task) related to the problem we are currently trying to solve (target task). In such cases, it would be wasteful to optimize the target task from scratch without utilizing the results of the source task. Therefore, it is expected that optimizing the target task can be made more efficient by transferring knowledge from the source task. This knowledge transfer is especially beneficial when evaluating the objective function is expensive, such as in hyperparamter optimization (Feurer and Hutter, 2019).
What is WS-CMA?.
Warm Starting CMA-ES (WS-CMA) (Nomura et al., 2021) accelerates the optimization of the target task by utilizing information from the source task. The process of WS-CMA is actually quite simple: Now we assume that there are solutions evaluated on the source task (Figure 8LABEL:sub@subfig:a_source_solutions). Based on the results of the source task, WS-CMA first estimates a promising distribution as Gaussian mixture models (GMM) (Figure 8LABEL:sub@subfig:b_promising_distribution). It then determines the parameters of the initial distribution to fit the promising distribution by minimizing the Kullback-Leibler divergence between GMM and the initial distribution (Figure 8LABEL:sub@subfig:c_initial_distribution). Subsequently, WS-CMA executes CMA-ES using the determined initial distribution to efficiently solve the target task.
How to use?.
Listing LABEL:list:ws-cma shows an example code using WS-CMA. We first need to prepare an array (source_solutions in the code) whose element is a tuple $(x,f(x))$. By using the method get_warm_start_mgd($\cdot$) with the created source_solutions, we can obtain the promising distribution parameters of the mean vector, the step-size, and the covariance matrix. WS-CMA is then performed by initializing CMA($\cdot$) with these parameters.
@@
4.3. Mixed-Integer Optimization
Why do we need CMAwM?.
While CMA-ES is basically performed for optimization of continuous variables, there are a variety of real-world problems that involve discrete variables such as binary or integers (Zhang et al., 2020; Miyagi et al., 2018; Feurer and Hutter, 2019). This raises the question: Can CMA-ES be applied to optimization problems involving a mix of continuous, binary, and integer variables?
A straightforward way to handle discrete variables in CMA-ES is to map continuous variables to discrete ones, i.e., encoding. For example, binary variables can be mapped as follows: $x_{i}(\in\mathbb{R})\leqslant 0.5$ is mapped to $0\in\{0,1\}$, and $x_{i}(\in\mathbb{R})>0.5$ to $1\in\{0,1\}$. However, this naïve approach may bias the search toward only 00s or $1$s, potentially leading to stagnation (Figure 9LABEL:sub@subfig:wo_margin).
What is CMAwM?.
Our library implements CMA-ES with Margin (CMAwM) (Hamano et al., 2022, 2024), a method developed for mixed-integer optimization and has been rapidly gaining attention recently (Watanabe et al., 2023; Thomaser et al., 2023; Andrei, 2024). To address the issue of the above naïve approach, CMAwM corrects distribution parameters to ensure that the probabilities of sampling 00 and $1$ remains above a certain threshold (Figure 9LABEL:sub@subfig:w_margin). This approach is also applicable to integer variables. By implementing this margin correction, CMAwM stably handles binary and integer variables while keeping the efficiency of CMA-ES for continuous variables.
Figure 9. Motivation for CMAwM. (a) Vanilla CMA-ES, which does not perform margin correction, risks biasing search toward only 00s or $1$s. (b) CMAwM introduces a lower bound on probability by margin correction, so that sample is not fixed to single discrete value.
To illustrate the impact of introducing the margin correction, we conducted experiments using CMA-ES with and without the margin. For simplicity, we considered a mixed-integer optimization problem, where the first $d_{\mathrm{co}}$ elements are continuous, and the remaining $d_{\mathrm{bi}}$ elements are binary. The encoding function is defined as follows:
(11) $\textsc{Enc}_{f}(x)=\left\{\begin{array}[]{lll}x_{i}&(1\leqslant i\leqslant d_{\mathrm{co}})\\ \mathbb{I}\{x_{i}>0.5\}&(d_{\mathrm{co}}+1\leqslant i\leqslant d_{\mathrm{co}}+d_{\mathrm{bi}})\end{array}\right.$
where $\mathbb{I}\{\cdot\}$ is the indicator function, and $d_{\rm co}$ and $d_{\rm bi}$ are the number of continuous and binary variables, respectively. The encoded solution is expressed as $\bar{x}=\textsc{Enc}_{f}(x)$. As an example, we employed the EllipsoidOneMax function which is defined as follows:
$f_{\rm EllipsoidOneMax}(\bar{x})=\sum_{j=1}^{d_{\mathrm{co}}}\left(1000^{\frac{j-1}{d_{\mathrm{co}}-1}}\bar{x}_{i}\right)^{2}+d_{\mathrm{bi}}-\sum_{k=d_{\mathrm{co}}+1}^{d_{\mathrm{co}}+d_{\mathrm{bi}}}\bar{x}_{k}.$
We set the number of dimension to $d_{\mathrm{co}}=d_{\mathrm{bi}}=20$. Figure 10 shows the results of the vanilla CMA-ES and CMAwM. We can observe that CMAwM successfully continues to improve the quality of the objective function value, whereas the vanilla CMA-ES has ceased to make such improvements.
Figure 10. Comparison between vanilla CMA-ES and CMAwM on the 40-D EllipsoidOneMax function. Initial distribution is $m^{(0}=[0.5,\ldots,0.5],\sigma^{(0)}=2.0,C^{(0)}=I$. Population size is set as $\lambda=4+\lfloor 3\log(d)\rfloor=15$. Vanilla CMA-ES stopped improving quality of objective function value, whereas CMAwM succeeds in finding optimum.
How to use?.
Listing LABEL:list:cmawm shows an example code using CMAwM on the 40-D EllipsoidOneMax function, which includes continuous and binary variables. The bounds variable defines the search space: np.tile([-np.inf, np.inf], ...) indicates that continuous variables are unconstrained; np.tile([0, 1], ...) is for binary variables. The steps variable specifies the discretization granularity: np.zeros is used for continuous variables and np.ones is used for binary variables. The ask method returns two values: x_eval and x_tell: x_eval is the encoded solution containing binary values, and is used for evaluation. x_tell is the raw solution that does not contain binary values, and is used for tell method. Our repository includes additional examples with integer variables.
@@ 4.3. Mixed-Integer Optimization
2from cmaes import CMAwM
@@
7bin_dim, cont_dim = 20, 20
8dim = bin_dim + cont_dim
9bounds = np.concatenate([
10 # for continuous variables
11 np.tile([-np.inf, np.inf], (cont_dim, 1)),
12 # for binary variables
13 np.tile([0, 1], (bin_dim, 1)),
14])
15# np.zeros is for continuous, np.ones is for binary
16steps = np.concatenate([np.zeros(cont_dim), np.ones(bin_dim)])
17# CMAwM needs bounds and steps for initialization
18optimizer = CMAwM(mean=np.zeros(dim), sigma=2.0, bounds=bounds, steps=steps)
19
20for _ in range(10000):
@@
23 x_eval, x_tell = optimizer.ask()
24 # x_eval is partially encoded in binary for eval.
25 value = ellipsoid_onemax(x_eval, binary_dim)
26 # x_tell is continuous and used for CMA update
27 solutions.append((x_tell, value))
@@
Listing 6: Example code of CMAwM on continuous and binary optimization problems.
5. Summary and Discussions
@@
This paper introduced cmaes, a simple yet practical CMA-ES library in Python. The library’s simplicity not only enhanced its software quality but also enabled easy integration with other libraries, such as Optuna (Akiba et al., 2019). Despite its simplicity, cmaes implemented highly practical methods that were recently proposed, such as LRA-CMA (Nomura et al., 2023), WS-CMA (Nomura et al., 2021), and CMAwM (Hamano et al., 2022), all with easy-to-use APIs. We hope that our library has become the ideal starting point for practitioners interested in exploring CMA-ES.
@@
It is important to acknowledge that the features in our library are not comprehensive. For example, pycma includes many features that our library lacks, as discussed in Section 1. Moreover, our library is not primarily designed for rigorous benchmarking tasks, such as comparing the performance of various optimizers. In this case, specialized software like COCO (Hansen et al., 2021) is recommended.

### v3

@@ 3.3. Unexpected Error Identification via Fuzzing
Fuzzing is an automated software testing method that aims to discover unexpected inputs that could cause program crashes. It involves generating a wide range of inputs. The key component in this process is the fuzzing engine, which automatically generates various inputs for the target program. A notable technique is coverage-guided fuzzing. This approach not only generates random inputs but also collects code coverage data. When an input is found to increase the code coverage, indicating untested paths, this information is used to methodically explore different branches within the program’s code. This strategy enhances the detection of unexpected errors, significantly contributing to the reliability of the software.
@@ 3.5. Integration with Real-World Systems:
A key factor in Optuna’s integration of cmaes is not only its clean code and simple design but also the reduced size of its pickle serialization. In hyperparameter optimization, the focus of Optuna, unexpected errors may occur during the training of machine learning models. Consequently, it is crucial to make the optimization process resumable, necessitating the regular saving of CMA-ES optimization data, including distribution parameters. Therefore, reducing the pickle size is critical for Optuna. Our library has introduced specialized __getstate__ and __setstate__ methods that are invoked during pickle serialization and deserialization. These modifications have successfully achieved further lightweighting of the pickle objects. Figure 3 shows the reduction of the serialization size for the CMA-ES objects using pickle in cmaes, compared to pycma.
Figure 3. Comparison of the serialization size for the CMA-ES objects after one generation using pickle in pycma and cmaes.
@@ 4. Examples with Recent Advances
  • LRA-CMA (Nomura et al., 2023): LRA-CMA can solve multimodal and noisy problems without expensive hyperparameter tuning.
  • WS-CMA (Nomura et al., 2021): WS-CMA can solve a current task efficiently by utilizing results of a similar task and setting good initial distributions (i.e. transfer learning).
  • CatCMAwM (Hamano et al., 2025): CatCMAwM can stably solve mixed-variable problems, including continuous, integer, and categorical variables.
  • COMO-CatCMAwM (Hamano et al., 2026): COMO-CatCMAwM uses multiple CatCMAwM instances to solve multi-objective mixed-variable problems.
Importantly, these methods are a generalization of CMA-ES and avoid unnecessary modifications to the original CMA-ES implementation. This enables the preservation of the library’s simplicity, making it easier for those who initially learned CMA-ES to straightforwardly delve into understanding these methods. Beyond these, advanced methods such as Safe CMA-ES (Uchida et al., 2024b) and CMA-ES on sets of points (CMA-ES-SoP) (Uchida et al., 2024a) have been implemented in cmaes.
@@ 4.2. Transfer Learning
Why do we need WS-CMA?
Practically it is possible that we have already solved a problem (source task) related to the problem we are currently trying to solve (target task). In such cases, it would be wasteful to optimize the target task from scratch without utilizing the results of the source task. Therefore, it is expected that optimizing the target task can be made more efficient by transferring knowledge from the source task. This knowledge transfer is especially beneficial when evaluating the objective function is expensive, such as in hyperparameter optimization (Feurer and Hutter, 2019).
What is WS-CMA?
Warm Starting CMA-ES (WS-CMA) (Nomura et al., 2021) accelerates the optimization of the target task by utilizing information from the source task. The process of WS-CMA is actually quite simple. First, we assume that there are solutions evaluated on the source task (Figure 6LABEL:sub@subfig:a_source_solutions). Based on the results of the source task, WS-CMA first estimates a promising distribution as Gaussian mixture models (GMM) (Figure 6LABEL:sub@subfig:b_promising_distribution). It then determines the parameters of the initial distribution to fit the promising distribution by minimizing the Kullback-Leibler divergence between GMM and the initial distribution (Figure 6LABEL:sub@subfig:c_initial_distribution). Subsequently, WS-CMA executes CMA-ES using the determined initial distribution to efficiently solve the target task.
How to use?
Listing LABEL:list:ws-cma shows an example code using WS-CMA. We first need to prepare an array (source_solutions in the code) whose elements are tuples $(x,f(x))$. By using the method get_warm_start_mgd($\cdot$) with the created source_solutions, we can obtain the promising distribution parameters of the mean vector, the step-size, and the covariance matrix. WS-CMA is then performed by initializing CMA($\cdot$) with these parameters.
@@
4.3. Mixed-Variable Optimization
Why do we need CatCMAwM?
Many real-world black-box optimization problems involve mixed decision variables defined over a search space composed of the continuous domain $\mathcal{X}$, the integer domain $\mathcal{Z}$, and the categorical domain $\mathcal{C}$. For example, hyperparameter optimization of deep learning models often requires jointly optimizing continuous variables (e.g., learning rates and regularization coefficients), integer variables (e.g., the number of layers), and categorical variables (e.g., choices of activation functions and optimization algorithms) (Feurer and Hutter, 2019; Elsken et al., 2019). Similar mixed-variable optimization problems are also common in domains such as materials design and chemical reaction optimization (Zhang et al., 2020; Ozaki et al., 2020a; Aldulaijan et al., 2024).
A straightforward method for handling discrete variables in CMA-ES is to map continuous samples to discrete values through rounding. However, as the multivariate Gaussian distribution converges, the generated discrete variables often become fixed to a (suboptimal) single value, leading to optimization stagnation. Furthermore, since categorical variables represent choices without an inherent numerical ordering, such discretization approaches impose unintended ordinal relationships that misrepresent the actual structure of the search space.
What is CatCMAwM?
To address these challenges, our library implements CatCMA with Margin (CatCMAwM), a stochastic optimization method designed for continuous, integer, and categorical variables. CatCMAwM is developed by incorporating novel integer handling into CatCMA (Hamano et al., 2024b), a mixed continuous-categorical black-box optimization method employing a joint distribution of multivariate Gaussian and categorical distributions. Although CatCMAwM targets mixed-variable optimization, its integer handling is more effective than the original one used in CMA-ES with Margin (CMA-ESwM) (Hamano et al., 2022, 2024a) even in mixed-integer optimization. This ensures a stable search by imposing both lower and upper bounds on the marginal probabilities , preventing premature fixation of discrete values while maintaining the convergence efficiency of continuous variables. On mixed-variable benchmarks, CatCMAwM achieves superior convergence performance compared to state-of-the-art Bayesian optimization methods.
How to use?
This CatCMAwM implementation provides a unified framework that automatically adapts to any combination of search spaces $\mathcal{X}$, $\mathcal{Z}$, and $\mathcal{C}$. As demonstrated in Listing LABEL:list:catcmawm, users can define the search space by specifying only the relevant arguments from x_space, z_space, and c_space; unused domains can be simply omitted, allowing the framework to automatically configure the search distribution. Figure 7 shows that the framework reduces to CMA-ES for continuous-only tasks, the Adaptive Stochastic Natural Gradient method (ASNG) (Akimoto et al., 2019) for categorical-only tasks, CMA-ESwM for mixed continuous-integer tasks, and CatCMA for mixed continuous-categorical tasks.
As shown in the loop of Listing LABEL:list:catcmawm, sampled solutions are returned as a Solution object, where specific variable sets are easily accessible via intuitive attributes such as sol.x, sol.z, and sol.c. This design allows users to focus solely on the relevant variable types for their problem while naturally ignoring unused domains.
@@ 4.3. Mixed-Variable Optimization
2from cmaes import CatCMAwM
@@
7# [lower_bound, upper_bound] for each continuous variable
8X = [[-5, 5], [-5, 5]]
9# possible values for each integer variable
10Z = [[-1, 0, 1], [-2, -1, 0, 1, 2]]
11# number of categories for each categorical variable
12C = [3, 3]
13
14optimizer = CatCMAwM(x_space=X, z_space=Z, c_space=C)
15
16for generation in range(50):
@@
19 sol = optimizer.ask()
20 value = SphereIntCOM(sol.x, sol.z, sol.c)
21 solutions.append((sol, value))
@@
Listing 6: Example code of CatCMAwM on continuous, integer, and categorical optimization problems.
Figure 7. By partially setting the variable domains, continuous domain x_space, integer domain z_space, and categorical domain c_space, users can run the appropriate algorithm within the CatCMAwM API without tedious case-by-case switching. The asterisk (*) indicates that CatCMAwM uses revised integer handling derived from CMA-ESwM.
4.4. Multi-Objective Optimization
Why do we need COMO-CatCMAwM?
Multi-objective optimization involves the simultaneous optimization of multiple, often conflicting, objective functions. Its primary goal is to identify a diverse set of Pareto-optimal solutions. For instance, multi-objective hyperparameter optimization (Morales-Hernández et al., 2022; Karl et al., 2023) targets objectives such as error-based performance, inference latency, and model size over the mixed-variable domain $\mathcal{X}\times\mathcal{Z}\times\mathcal{C}$. While multi-objective extensions of CMA-ES have been proposed for continuous variables (MO-CMA-ES (Igel et al., 2007a, b; Voß et al., 2010)) and mixed-integer variables (MO-CMA-ES with Margin (Hamano et al., 2024a)), these methods rely on elitist selection mechanisms, which are not straightforward to extend to categorical domains.
What is COMO-CatCMAwM?
COMO-CatCMA with Margin (COMO-CatCMAwM) is a multi-objective extension of CatCMAwM based on the Sofomore framework (Touré et al., 2019). This framework provides a general mechanism for constructing multi-objective optimizers by managing multiple single-objective algorithms, which are referred to as kernels. In COMO-CatCMAwM, each kernel is an instance of the CatCMAwM optimizer. As illustrated in Figure 8, these kernels are updated sequentially to optimize subproblems defined by the uncrowded hypervolume improvement (UHVI) fitness, utilizing the update rule of CatCMAwM. In bi-objective settings, COMO-CatCMAwM achieves competitive or superior hypervolume performance compared to established baselines such as NSGA-II (Deb et al., 2002) and MOTPE (Ozaki et al., 2020b, 2022).
Figure 8. Illustration of the Sofomore framework update in the objective space. The framework optimizes kernels sequentially using UHVI fitness, which integrates the hypervolume contribution and the distance to the empirical Pareto front (dotted line) with respect to the remaining incumbents.
How to use?
Listing LABEL:list:como-catcmawm shows an example code using COMO-CatCMAwM. Although COMO-CatCMAwM is technically applicable to problems with more than two objectives, our current implementation is limited to the bi-objective setting. Similar to the CatCMAwM API, users can define mixed-variable search spaces intuitively. Moreover, users do not need to track which internal kernel is currently being optimized and can simply use the tell($\cdot$) method, which automatically handles the sequential updates for the appropriate kernel.
1import numpy as np
2from cmaes import COMOCatCMAwM
3
4def DSIntLFTL(x, z, c):
5 ...
6 return [f1, f2]
7
8# [lower_bound, upper_bound] for each continuous variable
9X = [[-5, 15]] * 3
10# possible values for each integer variable
11Z = [range(-5, 16)] * 3
12# number of categories for each categorical variable
13C = [5] * 3
14
15optimizer = COMOCatCMAwM(x_space=X, z_space=Z, c_space=C)
16
17for _ in range(5000):
18 sol = optimizer.ask()
19 value = DSIntLFTL(sol.x, sol.z, sol.c)
20 optimizer.tell((sol, value))
Listing 7: Example code of COMO-CatCMAwM on bi-objective mixed-variable optimization problems.
Unlike other methods that use a fixed population size, the Sofomore framework requires a variable number of evaluations per step. This is because it evaluates both new candidates and current best solutions, called incumbents. To accommodate this variability, our library provides an iterator-based interface, ask_iter(), which yields all candidates currently requiring evaluation:
1solutions = []
2for sol in optimizer.ask_iter():
3 value = DSIntLFTL(sol.x, sol.z, sol.c)
4 solutions.append((sol, value))
5optimizer.tell(solutions)
As shown above, the tell($\cdot$) method can also receive a list of solution and evaluation value pairs. Notably, the order of submission to tell($\cdot$) is not required to match the sequence in which solutions were obtained from the ask interface. This flexibility allows the optimizer to handle results as they become available, making it highly compatible with asynchronous parallel evaluations.
5. Summary and Discussion
@@
This paper introduced cmaes, a simple yet practical CMA-ES library in Python. The library’s simplicity not only enhanced its software quality but also enabled easy integration with other libraries, such as Optuna (Akiba et al., 2019). Despite its simplicity, cmaes implements highly practical methods that were recently proposed, such as LRA-CMA (Nomura et al., 2023), WS-CMA (Nomura et al., 2021), CatCMAwM (Hamano et al., 2025), and COMO-CatCMAwM (Hamano et al., 2026), all with easy-to-use APIs. We hope that our library has become the ideal starting point for practitioners interested in exploring CMA-ES.
@@
It is important to acknowledge that the features in our library are not comprehensive. For example, pycma includes many features that our library lacks, as discussed in Section 1. Moreover, our library is not primarily designed for rigorous benchmarking tasks. In this case, specialized software like COCO (Hansen et al., 2021) is recommended. In addition, the current library, including the implementation of COMO-CatCMAwM, is limited to the bi-objective case, even though the method is technically applicable to problems with more than two objectives. Extending the implementation to general multi-objective settings is left for future work.